Lilian9667:

Graded Assignment: Volume Part 2 And I need help with it

Lilian9667:

1 attachment
Lilian9667:

I need help and supie said that your smart

AZ:

\(\text{Volume of a cone} = \pi r^2 \dfrac{h}{3}\)

Lilian9667:

Ok that makes sense Thank you

AZ:

so they told us radius = 6 and height = 8 can you put those numbers in place of r (for radius) and h (for height)

Lilian9667:

yes

Lilian9667:

Like on the picture?

AZ:

In the formula

Lilian9667:



AZ:

nice drawing! so let's plug it in \(\text{Volume of a cone} = \pi r^2 \dfrac{h}{3}\) \(\text{Volume of a cone} = \pi \times (6)^2 \times \dfrac{8}{3}\) can you multiply the numbers?

Lilian9667:

Thanks! Um the formula is V= 1/3 pi r2h

AZ:

that's correct

Lilian9667:

Yeah I do not know how to multiply that? lol

AZ:

okay, step by step what is 6^2 that's the same thing as 6 * 6 =

Lilian9667:

Oh ok

Lilian9667:

36

AZ:

so we go from V= 1/3 pi r^2 h V = 1/3 pi * (6)^2 * 8 V = 1/3 pi * 36 * 8 what is 36 * 8 =

Lilian9667:

288

AZ:

so now we have V = 1/3 pi * 36 * 8 V = 1/3 * pi * 288 now the first question wants an exact answer so that means we don't change pi to 3.14. We just leave it as pi so all we have left to do is 1/3 * 288 = ?

Lilian9667:

863

AZ:

uh oh careful there it's 1/3 not 3  so multiplying by 1/3 is the same as dividing by 3

Lilian9667:

So it's? =288 ---- 3

AZ:

yes

AZ:

what do you get when you divide

Lilian9667:

so divide 288 and 3

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ what do you get when you divide \(\color{#0cbb34}{\text{End of Quote}}\) good job - like always - congrats !!!

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 so divide 288 and 3 \(\color{#0cbb34}{\text{End of Quote}}\) yes

Lilian9667:

ok

Lilian9667:

96

AZ:

V = 1/3 * pi * 288 so we get our final answer for part A V = 96\(\pi\) for the second part, we just use 3.14 for pi so what is 96 * 3.14

Lilian9667:

It just keeps going Hehe

AZ:

yes which is why they asked you to just use 3.14

Lilian9667:

3.01.44

Lilian9667:

lol

AZ:

did you mean 301.44?

Lilian9667:

yah Thanks lol

AZ:

good job :)

Lilian9667:

Wait that was the answer?

Lilian9667:

V=301.44

AZ:

yes that's the answer to the second part

Lilian9667:

Thank you!

AZ:

You're welcome!